OpenStudy (anonymous):

A car averages 27 miles per gallon. If gas costs $4.04 per gallon, which of the following is closest to how much the gas would cost for this car to travel 2,727 typical miles? A. $ 44.44 B. $109.08 C. $118.80 D. $408.04 E. $444.40

OpenStudy (anonymous):

Since you're given the price of gas in gallons, what do you need to find? The number of gallens for the car to travel the 2727 miles right? How do we do this? First, lets find the number of gallons needed for 2727 miles. We can do this by dividing the miles by the miles per gallon, this way we have an number in gallons. So... 2727 / 27 = 101Gallons Now, we know we need 101 gallons, and since each gallon is 4.04, we can find how much it is, by multiplying the amount times the number needed So... 101 * 4.04 = 408.04 Your choice D is the correct answer. I know my answer is probably a bit longer than what you were looking for, but hopefully I wrote it out to the point that you'll understand the thought process and be able to do these problems in the future. Though, if you're lost on something just let me know!

OpenStudy (btaylor):

Write it out like this: \[ 2727 ~ miles \times \frac {1~ gallon}{27 ~miles} \times \frac{$4.04}{1 ~gallon}\]

OpenStudy (anonymous):

thank u so much @I_Need_Help_With_Home

OpenStudy (anonymous):

Your welcome @bluepizza